OpenStudy (anonymous):

Evaluate 6i^4 - 10i²

OpenStudy (anonymous):

\[i ^{2}=-1\] \[i ^{4}=(-1)^{2}=1\] can you finish it?

OpenStudy (anonymous):

6-10*(-1)=6+10=16

OpenStudy (anonymous):

is that 16i?

OpenStudy (anonymous):

-4 got ittttt thanks

OpenStudy (anonymous):

no ... the final answer should be 16.